### v16

# On the Bounds of Stopping and Cycle Numbers in the Collatz Process

Abstract.
We analyze the stopping-time and cycle structure of the normalized Collatz iteration. Using a recursive description of admissible binary sequences, we show that every integer $m\equiv 3\pmod{4}$ arises uniquely and derive new bounds for the associated stopping and cycle numbers. These bounds imply that $F_{q}(m)/m\to 1$ as the sequence length increases, while equality is impossible for any finite sequence. Consequently, no finite nontrivial cycle is compatible with the iteration, and the trivial cycle at $1$ is the only admissible periodic orbit.
Key words and phrases:
Collatz process, 3n+1 problem, Stopping time, Cycle number
2020 Mathematics Subject Classification:
Primary 11Y16, 00A05; Secondary 33D99, 68Q25
1. Introduction
The $3n+1$ problem, also known as the Collatz conjecture, has been widely studied and remains one of the most intriguing unsolved problems in mathematics [1], [2]. While several formulations of the Collatz function exist, we consider the following normalized form:
$f(n)=\begin{cases}F_{0}(n)=\frac{n}{2},&n=even\\ F_{1}(n)=\frac{3n+1}{2},&n=odd\end{cases}$
We denote the Collatz process as $F_{q}(n)=F_{101...0}(n)=F_{0}(...F_{1}(F_{0}(F_{1}(n))))$, where the binary sequence q=1010q=101...0, the total steps or the length of sequence is $s$ and the total number of odd steps is $r$. If $F_{q}(n)<n$, the minimum step $s$ is called the stopping time. If $F_{q}(m_{c})=m_{c}$, $m_{c}$ is called a cycle number. The stopping time ($s$) and the cycle numbers ($m_{c}$) have been widely studied in [3] - [9]. In [3], Simons and de Weger referred to a periodic sequence with $m$ local minima as an mcyclem-cycle. In the following discussion, we use the term ”cycle” to denote any sequence that contains at least one such cycle. We also restrict our analysis to odd starting values $n\equiv 3\pmod{4}$ satisfying $s\geq 4$, as those with $s\leq 2$ are readily confirmed by direct computation.
In this paper we obtain a general lower bound for the stopping number and compute sharp upper bounds for both stopping and cycle numbers across all admissible sequences. Taken together, these bounds show that any nontrivial cycle would require an exact identity between powers of 2 and 3 in the expression for $F_{q}(n)$, an equality unattainable for any finite sequence. Thus no finite nontrivial cycle can occur in the normalized Collatz iteration.
2. Collatz Sequence of Stopping
Let $q$ be a binary sequence of length $s$, containing exactly $r$ entries equal to 1. We define the function $1(i)$ to denote the position of the $i^{\text{th}}$ occurrence of the digit 1 within the sequence $q$. For example, consider $n=423$, $F_{q}(n)=302$, the sequence $q=1110110100$, which contains six ones. The corresponding positions of these ones are given by $1(1)=1$, $1(2)=2$, $1(3)=3$, $1(4)=5$, $1(5)=6$, and $1(6)=8$. Using this notation, the transformation $F_{q}(n)$ associated with the sequence $q$ in the Collatz process is given by
(2.1) $F_{q}(n)=\dfrac{3^{r}n+\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}},$
Here, $s$ is the total number of steps, $r$ is the number of applications of the function $F_{1}$, and $1(i)$ specifies the position of the $i^{\text{th}}$ occurrence of 1 in the sequence $q$.
Lemma 2.1.
Let $r\in\mathbb{N}$ and define
$s=\left\lfloor 1+r\cdot\log_{2}3\right\rfloor.$
Then there exist a non-negative integer $i\in\mathbb{Z}_{\geq 0}$ and a positive integer $k\in\mathbb{N}$ such that
$3^{r}\cdot(2i+1)-1=k\cdot 2^{s-r}.$
Proof.
By the definition of $s$, we have
$s=\left\lfloor 1+r\cdot\log_{2}3\right\rfloor\quad\Rightarrow\quad s-1<r\log_{2}3<s.$
Exponentiating both sides with base 2 gives
$2^{s-1}<3^{r}<2^{s},$
which implies
$\frac{3^{r}}{2^{s-r}}\in\left(2^{r-1},2^{r}\right).$
Now consider the congruence
$3^{r}\cdot t\equiv 1\pmod{2^{s-r}}$.
Since $3^{r}$ is odd and $\gcd(3^{r},2^{s-r})=1$, the inverse of $3^{r}$ modulo $2^{s-r}$ exists. Therefore, there exists an odd integer $t\equiv 2i+1$, for some $i\in\mathbb{Z}_{\geq 0}$, such that
$3^{r}\cdot(2i+1)\equiv 1\pmod{2^{s-r}}$.
Rewriting this congruence yields
$3^{r}\cdot(2i+1)-1=k\cdot 2^{s-r}$
for some $k\in\mathbb{N}$, as required. ∎
Lemma 2.2.
Let $r\geq 2,i\in\mathbb{N}$, and for any starting number
$m=2^{r}\cdot(2i+1)-1\equiv 3\pmod{4}$.
Then the Collatz sequence of $m$ begins with
$\underbrace{11\ldots 1}_{r\text{ times}}\,\underbrace{00\ldots 0}_{s-r\text{ times}}.$
Proof.
Let us begin with the expression
$m=2^{r}\cdot(2i+1)-1.$
We claim that applying $r$ iterations of the Collatz rule for odd integers, i.e., $m\mapsto(3m+1)/2$, leads to
$F_{11...1}(m)=3^{r}\cdot(2i+1)-1.$
This is shown by induction:
- Base case (step 1):
$\quad F_{1}(m)=\frac{3(2^{r}\cdot(2i+1)-1)+1}{2}=3\cdot 2^{r-1}(2i+1)-1=2^{r-1}\cdot 3^{1}\cdot(2i+1)-1.$
- Inductive step: Suppose that after r1r-1 steps,
$F_{11...1}(m)=2\cdot 3^{r-1}\cdot(2i+1)-1.$
Then, step $r$:
$F_{11...11}(m)=\frac{3[2\cdot 3^{r-1}\cdot(2i+1)-1]+1}{2}=\frac{2\cdot 3^{r}\cdot(2i+1)-2}{2}=3^{r}\cdot(2i+1)-1.$
So, after $r$ such ”odd steps” (each denoted by ‘1‘), we reach $3^{r}(2i+1)-1$, which is even (since $3^{r}$ is odd, so the result is even). From this point, the Collatz process proceeds with divisions by 2 only, i.e., steps of the form $m\mapsto m/2$, which we denote by ’0’.
Hence, the step sequence begins with exactly $r$ increasing steps (’1’s), followed by decreasing steps (’0’s) until the next odd number occurs.
Therefore, together with the Lemma 2.1, the sequence begins with:
$\underbrace{11\ldots 1}_{r\text{ times}}\,\underbrace{00\ldots 0}_{s-r\text{ times}},$
completing the proof. ∎
Table 1. Specific Sequences of $q$ with $1(i)=i$ for $m\equiv 3\pmod{4}$ in Collatz process
$r$ $s$ Sequence($q$) $2^{r}\cdot(2i+1)-1$ $m\pmod{2^{s}}$ $k=F_{q}(m)$
2 4 1100 $2^{2}\cdot 1-1$ 3 2
3 5 11100 $2^{3}\cdot 3-1$ 23 20
4 7 1111000 $2^{4}\cdot 1-1$ 15 10
5 8 11111000 $2^{5}\cdot 3-1$ 95 91
6 10 1111110000 $2^{6}\cdot 9-1$ 575 410
7 12 111111100000 $2^{7}\cdot 3-1$ 383 205
8 13 1111111100000 $2^{8}\cdot 1-1$ 255 205
9 15 111111111000000 $2^{9}\cdot 11-1$ 5,631 3,383
10 16 1111111111000000 $2^{10}\cdot 25-1$ 25,599 23,066
11 18 111111111110000000 $2^{11}\cdot 51-1$ 104,447 70,582
12 20 11111111111100000000 $2^{12}\cdot 17-1$ 69,631 35,291
13 21 111111111111100000000 $2^{13}\cdot 91-1$ 745,471 566,732
14 23 11111111111111000000000 $2^{14}\cdot 201-1$ 3,293,183 1,877,689
15 24 111111111111111000000000 $2^{15}\cdot 67-1$ 2,195,455 1,877,689
16 26 11111111111111110000000000 $2^{16}\cdot 193-1$ 12,648,447 8,113,298
17 27 111111111111111110000000000 $2^{17}\cdot 747-1$ 97,910,783 94,206,740
18 29 11111111111111111100000000000 $2^{18}\cdot 249-1$ 65,273,855 47,103,370
19 31 1111111111111111111000000000000 $2^{19}\cdot 83-1$ 43,515,903 23,551,685
20 32 11111111111111111111000000000000 $2^{20}\cdot 1393-1$ 1,460,666,367 1,185,813,152
21 34 1111111111111111111110000000000000 $2^{21}\cdot 3195-1$ 6,700,400,639 4,079,690,977
22 35 11111111111111111111110000000000000 $2^{22}\cdot 1065-1$ 4,466,933,759 4,079,690,977
23 37 1111111111111111111111100000000000000 $2^{23}\cdot 8547-1$ 71,697,432,575 49,111,434,902
24 39 111111111111111111111111000000000000000 $2^{24}\cdot 2849-1$ 47,798,288,383 24,555,717,451
25 40 1111111111111111111111111000000000000000 $2^{25}\cdot 22795-1$ 764,873,277,439 589,414,790,413
26 42 111111111111111111111111110000000000000000 $2^{26}\cdot 18521-1$ 1,242,923,270,143 718,351,699,928
27 43 1111111111111111111111111110000000000000000 $2^{27}\cdot 28019-1$ 3,760,646,520,831 3,260,217,528,257
28 45 111111111111111111111111111100000000000000000 $2^{28}\cdot 31185-1$ 8,371,159,695,359 5,442,907,506,622
29 46 1111111111111111111111111111100000000000000000 $2^{29}\cdot 10395-1$ 5,580,773,130,239 5,442,907,506,622
30 48 111111111111111111111111111111000000000000000000 $2^{30}\cdot 3465-1$ 3,720,515,420,159 2,721,453,753,311
31 50 11111111111111111111111111111110000000000000000000 $2^{31}\cdot 263299-1$ 565,430,297,034,751 310,197,425,018,629
Table 1 includes several specific sequences for values of $r$ ranging from 2 to 31. This list can be extended indefinitely, as there is no upper bound on $r$. Therefore, there is no upper bound on $s$, too. For example, when $r=1000$, then $s=1585$,
m=1131817111979028005238968810662181601027096670453939569503896194329480980139577293840980894314243553691246663949249137551627877867250646145606647381546299358516204398946163838768783595410569772717210118527009685388805997434584268801949287629806470254334768295217540415359597797337275682099623440815404654638275837876230714470651737163334210306119645953170309462667295881872408781238846687069393680025440093827021268188760756952836484710583975703867188025783792797611927080534015m=113181711197902800523896881066218160102709667045393956950389619432948098013957729384098\\ 0894314243553691246663949249137551627877867250646145606647381546299358516204398946163838768783\\ 5954105697727172101185270096853888059974345842688019492876298064702543347682952175404153595977\\ 9733727568209962344081540465463827583787623071447065173716333421030611964595317030946266729588\\ 1872408781238846687069393680025440093827021268188760756952836484710583975703867188025783792797\\ 611927080534015 (478 digits)
$F_{q}(m)=11027773766223939255599638243685383983123058040229118529895194635849895633361467411\\ 1382402477750726619658150157035536038452913603129669634101848918205329307515073658919130294253\\ 2781506097769243834069629788450644321665401676961180452393456153975938038355963675587842309211\\ 9591171602528480255231243448888268364113078890677027695508662577537129844756989406360976065975\\ 1963882667459405174174888685722485109676483971192668155890160044032425213449698721946740600735\\ 2136396783848666645$ (478 digits)
2i+1=1056283733409410554604530108264618017313129154566613687865030220907180023449545126574115508991430658543239682547832693931082637396808665045022380163533347206018898195558455646412i+1=105628373340941055460453010826461801731312915456661368786503022090718002344954512657\\ 411550899143065854323968254783269393108263739680866504502238016353334720601889819555845564641 (177 digits)
Theorem 2.3.
The summation $\sum_{i=1}^{r}3^{r-i}\cdot 2^{1(i)-1}\geq 3^{r}-2^{r},$ and equality holds when $1(i)=i$ for all $i$.
Proof.
To find the minimum value of $\sum_{i=1}^{r}3^{r-i}\cdot 2^{1(i)-1}$ where $1(i)$ is the position of the $i^{\text{th}}$ 1 in a binary sequence $q$ of length $s$, we apply the result from the Rearrangement Inequality. The minimum occurs when the positions $1(i)$ are as small as possible, i.e., $1(i)=i\quad\text{for }i=1,2,\dots,r$ That is, the 1s occur at the first $r$ positions in the sequence. Therefore, the minimum value is:
$\sum_{i=1}^{r}3^{r-i}\cdot 2^{i-1}$
We can compute this closed-form as a geometric convolution:
$\sum_{i=1}^{r}3^{r-i}\cdot 2^{i-1}=\sum_{j=0}^{r-1}3^{j}\cdot 2^{r-1-j}=2^{r-1}\sum_{j=0}^{r-1}\left(\frac{3}{2}\right)^{j}$
This is a geometric series:
$2^{r-1}\cdot\frac{\left(\frac{3}{2}\right)^{r}-1}{\frac{3}{2}-1}=2^{r-1}\cdot\frac{\left(\frac{3}{2}\right)^{r}-1}{\frac{1}{2}}=2^{r}\left[\left(\frac{3}{2}\right)^{r}-1\right]=3^{r}-2^{r}$
Final Answer:
$\boxed{\sum_{i=1}^{r}3^{r-i}\cdot 2^{1(i)-1}\geq 3^{r}-2^{r}}$
and equality holds when $1(i)=i$ for all $i$.
Theorem 2.4.
The summation
$\sum_{i=1}^{r}3^{\,r-i}\cdot 2^{\,1(i)-1}$
attains its maximum value when
$1(i)=\left\lfloor 1+(i-1)\log_{2}3\right\rfloor,\quad i=1,\dots,r.$
Proof.
Consider the modified rooted sequence tree introduced by Winkler [6], illustrated in Figure 1. It is evident that the further to the right the $1$’s appear in a sequence, the larger the resulting sum becomes comparing to the minimum summation when all $1$’s appear in the left. Consequently, the sequences along the top line of the sequence tree yield the maximum values.
Figure 1. Rooted sequence tree
Here is how these sequences in the modified sequence tree are constructed. All sequences are begin with $r=2,s=4$, so that the initial sequence is $q=1100$. This is ao called rooted tree. Each sub-sequence is obtained by inserting either “10” or “1” immediately before each zero at the terminal zero-block in the previous sequence, according to the difference in their lengths. Specifically, define
$d(r)=s(r)-s(r-1)=\left\lfloor 1+r\log_{2}3\right\rfloor-\left\lfloor 1+(r-1)\log_{2}3\right\rfloor.$
This rule generates all admissible sub-sequences. And the total number of the admissible sequences in each $r$ equal to the summary of all number of zeros at the terminal zero-block in the previous sequence. Consequently, the positions of the (1)’s in the top-line sequence are given exactly by
$1(i)=\left\lfloor 1+(i-1)\log_{2}3\right\rfloor,\quad i=1,\dots,r.$
This establishes the claim.
Table 2 shows the Collatz process of $r$ from 2 to 31, but with sequences with the maximum values of the summation. Again, the value of $r$ can be any big integers.
Table 2. Specific Sequences of $q$ with max value of the sum for $m\equiv 3\pmod{4}$ in Collatz process
$r$ $s$ Sequence($q$) $m\pmod{2^{s}}$ $F_{q}(m)<m$
2 4 1100 3 2
3 5 11010 11 10
4 7 1101100 59 38
5 8 11011010 123 118
6 10 1101101100 507 362
7 12 110110110100 1,019 545
8 13 1101101101010 3,067 2,458
9 15 110110110101100 7,163 4,304
10 16 1101101101011010 56,315 50,743
11 18 110110110101101100 23,547 15,914
12 20 11011011010110110100 416,763 211,226
13 21 110110110101101101010 941,051 715,420
14 23 11011011010110110101100 8,281,083 4,721,663
15 24 110110110101101101011010 4,086,779 3,495,268
16 26 11011011010110110101101100 62,807,035 40,287,332
17 27 110110110101101101011011010 96,361,467 92,716,039
18 29 11011011010110110101101101100 431,905,787 311,674,835
19 31 1101101101011011010110110110100 163,470,331 88,473,443
20 32 11011011010110110101101101101010 1,237,212,155 1,004,406,265
21 34 1101101101011011010110110110101100 7,679,663,099 4,675,937,150
22 35 11011011010110110101101101101011010 33,449,466,875 30,549,700,432
23 37 1101101101011011010110110110101101100 119,348,812,795 81,751,762,091
24 39 110110110101101101011011011010110110100 462,946,196,475 237,832,281,869
25 40 1101101101011011010110110110101101101010 188,068,289,531 144,926,270,443
26 42 110110110101101101011011011010110110101100 2,936,847,358,971 1,697,360,845,538
27 43 1101101101011011010110110110101101101011010 5,135,870,614,523 4,452,440,639,554
28 45 110110110101101101011011011010110110101101100 18,330,010,147,835 11,918,127,650,276
29 46 1101101101011011010110110110101101101011011010 35,922,196,192,251 35,034,785,816,635
30 48 110110110101101101011011011010110110101101101100 211,844,056,636,411 154,958,046,921,632
31 50 11011011010110110101101101101011011010110110110100 915,531,498,413,051 502,264,407,868,691
Lemma 2.5.
For any integer $s\geq 1$,
$\sum_{i=0}^{s-1}2^{\,i}=2^{\,s}-1.$
Proof.
Consider the finite geometric series with ratio $2$:
$\sum_{i=0}^{s-1}2^{\,i}=\frac{2^{\,s}-1}{2-1}=2^{\,s}-1.$
Lemma 2.6.
Let $m\equiv 3\pmod{4}$. Then $m$ admits a unique representation of the form
$m=3+\sum_{n_{i}\geq 2}2^{n_{i}},\qquad n_{i}\geq 2,$
where the exponents $n_{i}$ are distinct. Conversely, every finite sum of the form
$3+\sum_{n_{i}\geq 2}2^{n_{i}},\qquad n_{i}\geq 2,$
yields an integer $m\equiv 3\pmod{4}$.
Proof.
Since $2^{n}\equiv 0\pmod{4}$ for all $n\geq 2$, any integer of the form $3+\sum_{n_{i}}2^{n_{i}}$ is congruent to $3\pmod{4}$. Conversely, if $m\equiv 3\pmod{4}$, then m3m-3 is divisible by 4 and hence has a binary expansion involving only powers $2^{n}$ with $n\geq 2$. Uniqueness follows from the uniqueness of the binary expansion.
Example 2.7.
For instance,
$\begin{array}[]{c|l}4i-1\equiv 3\pmod{4}&\sum_{n_{i}\geq 2}2^{n_{i}}+3\\ \hline\cr 3&3\\ 7&2^{2}+3\\ 11&2^{3}+3\\ 15&2^{3}+2^{2}+3\\ 19&2^{4}+3\\ 23&2^{4}+2^{2}+3\\ 27&2^{4}+2^{3}+3\\ 31&2^{4}+2^{3}+2^{2}+3\\ 35&2^{5}+3\\ 39&2^{5}+2^{2}+3\\ 43&2^{5}+2^{3}+3\\ 47&2^{5}+2^{3}+2^{2}+3\\ 51&2^{5}+2^{4}+3\\ 55&2^{5}+2^{4}+2^{2}+3\\ 59&2^{5}+2^{4}+2^{3}+3\\ 63&2^{5}+2^{4}+2^{3}+2^{2}+3\\ 67&2^{6}+3\\ 71&2^{6}+2^{2}+3\\ 75&2^{6}+2^{3}+3\\ 79&2^{6}+2^{3}+2^{2}+3\\ 83&2^{6}+2^{4}+3\\ 87&2^{6}+2^{4}+2^{2}+3\\ 91&2^{6}+2^{4}+2^{3}+3\\ 95&2^{6}+2^{4}+2^{3}+2^{2}+3\\$ $\vdots$$\vdots\end{array}$
Lemma 2.8 (Uniqueness of sequences and starting values).
The binary sequences $q$ displayed in Figure 1 are pairwise distinct. Moreover, the construction defines an injective mapping
$q\;\longmapsto\;m,$
where each sequence $q$ corresponds to a unique integer $m\equiv 3\pmod{4}$ satisfying $F_{q}(m)<m$.
Proof.
The sequences in Figure 1 are generated recursively using a deterministic insertion rule: at each stage $r$, a new sequence is obtained from a unique parent sequence by inserting either the substring $1$ or $10$ at a prescribed position immediately preceding the terminal block of zeros. Since both the parent sequence and the insertion position are uniquely determined, distinct construction paths necessarily produce distinct binary sequences. Hence all sequences $q$ in Figure 1 are pairwise distinct.
For a fixed sequence $q$ of length $s$, equation (2.1) uniquely determines the residue class of $m\pmod{2^{s}}$. By Lemma 2.6, every integer $m\equiv 3\pmod{4}$ admits a unique binary expansion of the form
$m=3+\sum_{n_{i}\geq 2}2^{n_{i}}.$
The exponents $n_{i}$ correspond uniquely to the positions of the $1$’s in the sequence $q$. Consequently, distinct sequences $q$ give rise to distinct integers $m$. Finally, direct comparison of the iterates of $m$ with its binary expansion shows that the constructed integer satisfies $F_{q}(m)<m$. This establishes injectivity of the mapping $q\mapsto m$. ∎
Example 2.9.
For instance, let $i$ is any integer and $4i-1\equiv 3\pmod{4}$, then the first two iterates are
$F_{1}(4i-1)=3\cdot 2i-1,\qquad F_{11}(4i-1)=3^{2}i-1.$
As an illustrative example, take
$r=11,\quad s=18,\quad q=111111111010100000.$
Then the iterates proceed as
$F_{111}(4i-1)$ $=\frac{3(3^{2}(2j)-1)+1}{2}=3^{3}j-1,$
$F_{1111}(4i-1)$ $=\frac{3(3^{3}(2k)-1)+1}{2}=3^{4}k-1,$
$F_{11111}(4i-1)$ $=\frac{3(3^{4}(2\ell)-1)+1}{2}=3^{5}\ell-1,$
$F_{111111}(4i-1)$ $=\frac{3(3^{5}(2m)-1)+1}{2}=3^{6}m-1,$
$F_{1111111}(4i-1)$ $=\frac{3(3^{6}(2n)-1)+1}{2}=3^{7}n-1,$
$F_{11111111}(4i-1)$ $=\frac{3(3^{7}(2o)-1)+1}{2}=3^{8}o-1,$
$F_{111111111}(4i-1)$ $=\frac{3(3^{8}(2p)-1)+1}{2}=3^{9}p-1,$
[…]
The corresponding expansion of 4i14i-1 in powers of two is
$4i-1$ $=2^{2}(2j)-1=2^{2}2^{2}k-1=2^{2}2^{3}\ell-1=2^{2}2^{4}m-1=2^{2}2^{5}n-1=2^{2}2^{6}o-1$
[…]
$=2^{s}y+\sum_{n_{i}\geq 2}2^{n_{i}}+3=2^{s}y+m.$
[…]
The construction in Figure 1 generates each integer $m\equiv 3\pmod{4}$ exactly once. Equivalently, the correspondence
$q\;\longleftrightarrow\;m$
defined by Figure 1 is a bijection between admissible sequences $q$ and integers $m\equiv 3\pmod{4}$.
Proof.
We prove the statement by showing that the constructions $q\mapsto m$ and $m\mapsto q$ are well-defined and inverse to each other.
Forward direction ($q\mapsto m$). Starting from the root $q=11$, we have
$F_{11}(4i-1)=3^{2}i-1.$
Any admissible sequence $q=11\cdots$ corresponds to a path in the rooted sequence tree shown in Figure 1. Along this path, the Collatz process, $F_{q}(4i-1)=3^{r}y+F_{q}(m)$, yields successive decompositions
[…]
Consequently, we may write
$4i-1=2^{s}y+\sum_{n_{i}\geq 2}2^{n_{i}}+3=2^{s}y+m,$
which implies
$m\equiv 3\pmod{4}$.
Backward direction ($m\mapsto q$). Let $m$ be any integer satisfying $m\equiv 3\pmod{4}$. By Lemma 2.6, $m$ admits a representation of the form
$m=\sum_{n_{i}\geq 2}2^{n_{i}}+3.$
Writing $4i-1=2^{r}y+m$ and using the powers $2^{n_{i}}$ in the above decomposition, we recover successive relations
$i=2j\text{ or }2j+1,\quad j=2k\text{ or }2k+1,\quad k=2l\text{ or }2l+1,\ \ldots,$
which uniquely determine an admissible sequence
$q=11\cdots$
can be obtained via the inverse Collatz process and the mapping $F_{q}(4i-1)=3^{r}y+F_{q}(m)$.
Since the mappings $q\mapsto m$ and $m\mapsto q$ are inverses of each other, the correspondence between admissible sequences $q$ and integers $m\equiv 3\pmod{4}$ is bijective. ∎
Theorem 2.11.
For any integer $i$, we have
$4i-1\equiv 3\pmod{4},\quad\text{and}\quad F_{q}(4i-1)\neq 4i-1.$
Proof.
From above, we have
$4i-1=2^{s}y+m,$
and
$F_{q}(4i-1)=3^{r}y+F_{q}(m),$
where $y$ is any integer.
Suppose, for the sake of contradiction, that
$4i-1=F_{q}(4i-1).$
Then it follows that
$m=F_{q}(m),$
and hence
$2^{s}y=3^{r}y.$
Since $y\neq 0$, we obtain
$2^{s}=3^{r}.$
This is impossible, because no power of $2$ can equal a power of $3$. Therefore,
$F_{q}(4i-1)\neq 4i-1,$
which completes the proof. ∎
Remark: This implies that for any finite sequence, the existence of a loop is impossible.
3. Bounds of Stopping Number and Cycle Number
Suppose that there exists a cycle in the sequence, i.e.
$F_{q}(m_{c})=m_{c}.$
Using equation (2.1), we have
$\dfrac{3^{r}m_{c}+\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}}=m_{c}.$
Thus,
$m_{c}=\dfrac{\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}-3^{r}}.$
Theorem 3.1.
The non-trivial cycle number $m_{c}$ is lower bounded by
$\boxed{m_{c}\geq\dfrac{3^{r}-2^{r}}{2^{s}-3^{r}}}$
Proof.
Using Theorem 2.3, we immediately obtain
$m_{c}=\dfrac{\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}-3^{r}}\geq\min_{q}\dfrac{\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}-3^{r}}=\dfrac{3^{r}-2^{r}}{2^{s}-3^{r}},$
as claimed.
Remark: Under the condition of Lemma 2.1, i.e., $s=\left\lfloor 1+r\cdot\log_{2}3\right\rfloor$, thus $3^{r}<2^{s}$ holds, which makes the denominator $2^{s}-3^{r}$ positive.
For example, the value of expression
$\frac{3^{r}-2^{r}}{2^{s}-3^{r}}\quad\text{for }r=431{,}166{,}034{,}846{,}567\text{ and }s=683{,}381{,}996{,}816{,}440$
is approximately: 7.50×1014. The values of $r$ and $s$ used here correspond to the maximum values listed in the reference [7]. Based on the results from [8], the upper bound for the number of non-trivial cycles can be expressed as
$\boxed{m_{c}<\dfrac{s\cdot 3^{r}}{2^{s}-3^{r}}.}$
Substituting r=431,166,034,846,567r=431{,}166{,}034{,}846{,}567 and s=683,381,996,816,440s=683{,}381{,}996{,}816{,}440, this upper bound evaluates to approximately 5.12×1029.
Rewrite the stopping number as,
$F_{q}(m)=\dfrac{3^{r}m+\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}}=\dfrac{3^{r}}{2^{s}}m+\sigma_{q},$
where
$\sigma_{q}=\dfrac{\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}},$
it is easy to see that the stopping number $F_{q}(m)$ consists of two terms, the first term includes starting number $m$ and is unrelated to the sequence $q$, while the second term doesn’t include starting number $m$, but it is only related to the sequence. The value of $F_{q}(m)$ is dominated by the first term. The second term determines the position of $F_{q}(m)$ between $m$ and $m/2$. Similar to the the cycle number, the lower bound of the second term $\sigma_{q}$ is determined by
$\boxed{\sigma_{q}\geq\dfrac{3^{r}-2^{r}}{2^{s}},}$
while the upper bound based on the results of [8] is
$\boxed{\sigma_{q}<\dfrac{s3^{r}}{2^{s}},}$
and from the observation data that are showing in Figure 2. Here, the starting numbers $m$ for Figure 2A, B, and C are from the positive integer of $3\pmod{4}$ start from 3 to 16,000,000 whose stopping time $s\geq 4$. We can see that
$\boxed{\sigma_{q}<\alpha\dfrac{3^{r}-2^{r}}{2^{s}},}$
where $\alpha$ is a constant dependent on the maximum value of $r$ in all sequences.
Figure 2. Collatz process results for integers $m\equiv 3\pmod{4}$, ranging from 3 to 16,000,000. A. Second term $\sigma_{q}$ (blue dots) and its bounds in green and red versus $3^{r}/2^{s}$; B. Ratio $F_{q}(m)/m$ (blue dots) versus $3^{r}/2^{s}$; C. Starting number $m$ (blue dots), cycle number $m_{c}$ in green and its bounds in red and black versus $3^{r}/2^{s}$; D. Estimated bounds for the cycle number $m_{c}$ versus $3^{r}/2^{s}$ using extended $r$.
From the computing results, we observed that α=max(r)/4. Here $\alpha$=10 for Figure 2A,B and C; $\alpha$=160 for Figure 2D with max of $r=640$. We also note that the estimated bond by Luca (2005) for cycle numbers in [8] has the same pattern as our calculated max of cycle numbers that plotted in the Figure 2D. The relationship between them can be regressed as y=6.58x2.37, which is plotted in the Figure 3.
Figure 3. Regression of Estimated bound for the cycle number $m_{c}$ versus Calculated max values with $max(r)=640$.
Figure 2A shows the second term of $F_{q}(m)$ (i.e. $\sigma_{q}$) and its bounds versus the ratio of $3^{r}$ and $2^{s}$. The Figure 2B shows that the non-trivial cycle number would be where $\dfrac{3^{r}}{2^{s}}$ is close to 1 if it existed. Figure 2B shows the ratio $F_{q}(m)/m$ versus ratio of $3^{r}$ and $2^{s}$. Figure 2C shows the starting number $m$ (blue dots) and the cycle number $m_{c}$ (green line) and its bounds (red lines). Figure 2D is showing the bounds of the cycle number with more values of $3^{r}/2^{s}$. The data in this figure was generated for $r=1$ to 640 (645 is the maximum number that R software can go with), then we apply $s=\left\lfloor 1+r\cdot\log_{2}3\right\rfloor$. The value of $s$ corresponding to $r$ is unique because of this restriction. Then we use the pair of $(r,s)$ to calculate the cycle number bound. We found that the minimum cycle number $m_{c}$ is greater than 978 when $(r,s)=(306,485)$ and $3^{r}/2^{s}\approx 0.9989783$. It is not hard to see that the ratio $F_{q}(m)/m$ tends to be 1 as ratio $3^{r}/2^{s}$ goes to 1. Therefore, if there exists a cycle, the inequality
$\dfrac{3^{r}}{2^{s}}+\alpha\dfrac{3^{r}-2^{r}}{2^{s}m}\geq 1$
must be hold as $3^{r}/2^{s}$ tends to be 1. Thus, the cycle number $m_{c}$ must be less than a certain number, i.e.
(3.1) $\boxed{m_{c}\leq\alpha\dfrac{3^{r}-2^{r}}{2^{s}-3^{r}},}$
which is also can be validate by the computing results in Figure 2C.
Finally, we have the following theorem.
Theorem 3.2.
The inequalities
$1-\epsilon_{1}=\dfrac{1}{\alpha}+(1-\dfrac{1}{\alpha})\dfrac{3^{r}}{2^{s}}<\frac{F_{q}(m)}{m}<\alpha+(1-\alpha)\dfrac{3^{r}}{2^{s}}=1+\epsilon_{2}$
hold, and $\lim_{r\to\infty}\tfrac{F_{q}(m)}{m}\to 1$.
Remark 3.3.
This shows that any nontrivial cycle, if it exists, must be extraordinarily large and finely tuned, consistent with all known computational and theoretical bounds.
Proof.
Building on the above discussion, and applying Theorem 2.3 along with the established lower bound Theorem 3.1 and upper bounds Equation 3.1 for the cycle number, we obtain the following results:
$\dfrac{F_{q}(m)}{m}$ $=\dfrac{3^{r}}{2^{s}}+\dfrac{\sum_{i=1}^{r}3^{r-i}2^{1(i)-1}}{2^{s}m}$
$\geq\dfrac{3^{r}}{2^{s}}+\dfrac{3^{r}-2^{r}}{2^{s}m}$
$>\dfrac{3^{r}}{2^{s}}+\dfrac{2^{s}-3^{r}}{2^{s}\alpha}$
$=\dfrac{1}{\alpha}+(1-\dfrac{1}{\alpha})\dfrac{3^{r}}{2^{s}}$
$=1-\epsilon_{1},$
and
$\dfrac{F_{q}(m)}{m}$ $<\dfrac{3^{r}}{2^{s}}+\alpha\dfrac{3^{r}-2^{r}}{2^{s}m}$
$\leq\dfrac{3^{r}}{2^{s}}+\alpha\dfrac{2^{s}-3^{r}}{2^{s}}$
$=\alpha+(1-\alpha)\dfrac{3^{r}}{2^{s}}$
$=1+\epsilon_{2},$
respectively, so that
$1-\epsilon_{1}=\dfrac{1}{\alpha}+(1-\dfrac{1}{\alpha})\dfrac{3^{r}}{2^{s}}<\frac{F_{q}(m)}{m}<\alpha+(1-\alpha)\dfrac{3^{r}}{2^{s}}=1+\epsilon_{2}.$
Given $\alpha=max(r)/4$, then we have $\lim_{r\to\infty}\alpha\to\infty$, and we know that $\lim_{r\to\infty}\dfrac{3^{r}}{2^{s}}\to 1$, therefore, $\lim_{r\to\infty}\tfrac{F_{q}(m)}{m}\to 1$.
The value of $1-\dfrac{3^{r}}{2^{s}}$ for $r=431{,}166{,}034{,}846{,}567\text{ and }s=683{,}381{,}996{,}816{,}440$, is approximately: 1.33×1015.
The Corollary 1.8 in [10] gives that
$1-\dfrac{3^{r}}{2^{s}}\geq(e\cdot s)^{-C},$
where $C=e\cdot 2^{3.5}30^{5}\log 3\approx 821013301.$ Thus,
$(e\times 683{,}381{,}996{,}816{,}440)^{-821{,}013{,}301}\approx\boxed{10^{-7.1\times 10^{9}}}.$
Although this is unimaginably small—essentially zero for any computational purpose, $1-\dfrac{3^{r}}{2^{s}}\neq 0$, as $3^{r}\neq 2^{s}$.
4. Conclusion
The results obtained in this paper indicate that the stopping time in the Collatz process admits no finite upper bound. For admissible sequences $q$ with parameters $(r,s)$, the ratio $3^{r}/2^{s}$ tends to $1$ as the sequence length increases, yet it never equals $1$. Consequently, the quotient $F_{q}(m)/m$ approaches but never attains unity. This structural obstruction rules out the possibility of a finite nontrivial cycle, since any such cycle would require an exact equality between powers of $2$ and $3$. Combined with the lower and upper bounds established in this work, these observations reinforce the conclusion that all nontrivial cycles are excluded, and that the trivial cycle at $1$ is the only admissible periodic behavior.
Supplemental Materials: R Programs
Below is the R code used for generating all numbers $s,m,F_{q}(m),$ sequences $q$ in Figure 1, and many others based on $r$.
# function of even integer
is.even <- function(x) {
if (as.bigz(x) %% 2 == 0) TRUE
else FALSE
# function to insert a substring in a string at certain position
insert_substr <- function(orig_str, substr, pos) {
# Split the original string into two parts
part1 <- substr(orig_str, 1, pos - 1)
part2 <- substr(orig_str, pos, nchar(orig_str))
# Concatenate the parts with the substring in between
result <- paste0(part1, substr, part2)
return(result)
# initialize
e<-NULL
n<- NULL
E<- NULL
r=2
s=floor(1+r*log23)
q <- "1100"
k=1
e[k]=2 # # of zeros of the sequence tail
E[[r]]\$e<-e
E[[r]]\$q<-q
# not necessary
E[[r]]\$m<-3
E[[r]]\$d<-2
E[[r]]\$r<-2
E[[r]]\$s<-4
E[[1]]\$r<-1
E[[1]]\$s<-2
E[[1]]\$e<-1
E[[1]]\$q<-’10’
E[[1]]\$m<-1
E[[1]]\$d<-1
mk<-NULL
Fmk<-NULL
# start from 3:29 took 3.698423559202088206632 hours
for (r in 3:12){
s=floor(1+r*log23)
s1=floor(1+(r-1)*log23)
d=s-s1
ins <- ifelse(d == 2, "10", "1")
pre_q <- E[[r-1]]\$q
e1<-E[[r-1]]\$e
n <- sum(e) # No. of residue classes in m(mod2^s)
k=1
# for each previous q’s’
for (u in 1:length(e1)){
j=d
q0<-pre_q[u]
# get new q’s
for (i in 1:e1[u]) {
position <- s1-i+1
q[k] <- insert_substr(q0, ins, position)
[…]
e[k] <- j
j=j+1
# get m and Fm from q created above
# initialize the values and integer type (big z)
m=-1
Fm=-1
p=2
t=2
# use when s>50, about 3X slow
# m=-big_1
# Fm=-big_1
# p=big_2
# t=big_2
for (i in 3:s) {
step <- substr(q[k], i, i)
if (step==0) {
p=p #power of 3, max=r
if (is.even(Fm)) {
Fm=Fm/2
m=m
} else {
Fm=(3^p+Fm)/2
m=2^t+m
} else if (step==1) {
p=p+1
if (is.even(Fm)) {
Fm=(3*(3^(p-1)+Fm)+1)/2
m=2^t+m
} else {
Fm=(3*Fm+1)/2
m=m
t=t+1 # power of 2 max=s
mk[k]<-m
Fmk[k]<-Fm
# Using big number package
# mk[k]<-as.numeric(m)
# Fmk[k]<-as.numeric(Fm)
k=k+1
E[[r]]<-list(r=r,s=s,d=d,e=e,n=n,q=q, m=mk,Fm=Fmk)